Can you please explain step by step on how to do each of these two problems below? And then tell me whether it involves an arithmetic sequence or a geometric sequence I also need the work shown out for both problems
thank you.

Perform the indicated operations and reduce the answer to lowest terms.
(-3/4) ÷ (5/8)
(9/14 ÷ 3/7) × ½

Answer by cleomenius(959)

Divide fractions by inverting the numerator and denominator, and multiplying.
-3/4 * 8/5 = -24/20 = -6/5
9/16 * 7/3 = 63/48 * 1/2 = 63/96
I don't believe either problem at this point would be involve either a geometric or arithmetic sequence.
Cleomenius.